OpenStudy (anonymous):

Find y' when y=e^sinhx cosh (3x)

OpenStudy (nikita2):

sinhx = (e^x - e^-x)/2. (sinhx)' = (e^x + e^-x)/2 = cosh(x). (cosh(x))' = (e^x - e^-x)/2 = sinhx. So y' = e^sinhx cosh (3x)*( coshx*cosh3x + sinhx*3sinh3x ).

OpenStudy (anonymous):

wahooooooooooooooooooooo, thanks :)